kekeman:

A large emerald with a mass of 982.7 grams was recently discovered in a mine. If the density of the emerald is 2.76grams over centimeters cubed, what is the volume? Round to the nearest hundredth when necessary, and only enter numerical values, which can include a decimal point.

sophia333:

Density * Volume = Mass. Correct?

sophia333:

M = D/V.

sophia333:

Shortening it to variables ^^

sophia333:

D= 2.76. As stated in the paragraph.

sophia333:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @sophia333 M = D/V. \(\color{#0cbb34}{\text{End of Quote}}\) V = M/D Sorry

sophia333:

M=982.7

sophia333:

Divide those together, what is your answer?

sophia333:

@kekeman

sophia333:

You can use your QuestionCove Calculator.

kekeman:

356.05?

kekeman:

Then what is the answer

kekeman:

I need proof that this is right

kekeman:

Because i got 356.05

sophia333:

Hmm well lemme show u

sophia333:

Oh wait wait sorry you are correct, i did the problem incorrectly.

sophia333:

I apologize.

kekeman:

All good